OpenStudy (anonymous):

6/2(1+2) = ? I'd say 9 but I'm not 100% certain.

OpenStudy (lgbasallote):

yes it's 8

OpenStudy (lgbasallote):

9*

OpenStudy (anonymous):

8!?

OpenStudy (anonymous):

oh :p

OpenStudy (lgbasallote):

6/2 = 3 then 3(1 + 2) = 3(3) = 9

OpenStudy (anonymous):

right ok. I wasn't sure if I was supposed to multiply with the brackets first

hartnn (hartnn):

1

Parth (parthkohli):

\[{6 \over 2}(1 + 2) = {6 \over 2}(3) = 3(3) = 9\]

hartnn (hartnn):

my calculator gave 1

Parth (parthkohli):

Although, the answer is 1 when the question is the following:\[{6 \over 2(1 + 2)}\]But 9 is the correct answer through PEMDAS.

OpenStudy (lgbasallote):

nope @xcrypt according to PEMDAS rule...when you have multiplication and division, you go from left to right

OpenStudy (anonymous):

yes true.

Parth (parthkohli):

You must always use the rules of PEMDAS. Parentheses: 6/2(3) Exponent(none) Multiplication and division: left to right - 3(3) Multiply 3 and 3: 9